# Markovian dynamics for a quantum/classical system and quantum trajectories

### v1

@@ Abstract
Quantum trajectory techniques have been used in the theory of open systems as a starting point for numerical computations and to describe the monitoring of a quantum system in continuous time. Here we extend this technique and use it to develop a general approach to the dynamics of quantum/classical hybrid systems. By using two coupled stochastic differential equations, we can describe a classical component and a quantum one which have their own intrinsic dynamics and which interact with each other. A mathematically rigorous construction is given, under the restriction of having a Markovian joint dynamics and of involving only bounded operators on the Hilbert space of the quantum component. An important feature is that, if the interaction allows for a flow of information from the quantum component to the classical one, necessarily the dynamics is dissipative. We show also how this theory is connected to a suitable hybrid dynamical semigroup, which reduces to a quantum dynamical semigroup in the purely quantum case and includes Liouville and Kolmogorov-Fokker-Plank equations in the purely classical case. Moreover, this semigroup allows to compare the proposed stochastic dynamics with various other proposals based on hybrid master equations. Some simple example are constructed in order to show the variety of physical behaviours which can be described; in particular, a model presenting hidden entanglement is introduced.
@@ 1 Introduction
The interest in quantum/classical hybrid systems is old, see for instance [1, 7, 14, 13, 4, 3, 8, 6, 10, 11, 9, 12, 2, 5] and references therein. One of the main motivations of the study of hybrid systems is that the output of a monitored system is classical; then, implicitly or explicitly, the dynamics of quantum/classical systems is involved in the theory of quantum measurements in continuous time, quantum filtering,…[15, 19, 23, 25, 1, 2, 18, 22, 20, 24, 16, 17, 21, 4, 3]. Moreover, hybrid systems could be used as an approximation to complicate quantum systems, as an effective theory [1, 4, 8, 9, 12, 11]. Another motivation is in the connections between a theory of gravity and quantum theories; perhaps gravity is “classical”, and a dynamical theory of gravity and matter needs a hybrid dynamics [28, 5, 26, 4, 1, 27]. Classical environments could be involved even in revivals of quantum entanglement [29, 30]. A common feature in many of these attempts is that a non-trivial hybrid dynamics is always irreversible, due to the flow of information from the quantum system to the classical one [1, 6, 4, 26, 14, 13, 7].
The main aim of this article is to construct a consistent formulation of the dynamics of quantum/classical hybrid systems, by connecting the quantum trajectory theory, used to describe open and monitored quantum systems [31, 32, 23, 24, 16, 17, 21, 23], to the classical theory of Markov processes and stochastic differential equations [33, 34, 35, 36, 37, 39, 38]. The dynamics of the hybrid systems is given in Sec. 2 in terms of coupled stochastic differential equations; only the Markovian case is considered. We also show how this approach is connected to notions appearing in the general formulation of quantum measurements, such as “positive operator valued measures” and “instruments” [24, 22].
In Sec. 3 we show how to associate a suitable semigroup to the constructed dynamics. This hybrid semigroup reduces to a quantum dynamical semigroup [22, 24] in the case of a pure quantum system, while in the pure classical case it includes the Liouville and the Kolmogorov-Fokker-Planck equations [39]. We give also the formal generator of this semigroup, which allows to compare the class of dynamics developed here with the ones obtained by other approaches. This generator also shows analogies and links with the theory of infinitely divisible distributions [39, 40, 41, 13].
@@ 1.1 Some notations
By $\mathscr{S}(\mathscr{H})\subset\mathscr{T}(\mathscr{H})$ margin: $\mathscr{S}(\mathscr{H})$ we denote the set of the statistical operators (the quantum states). Then, general measurements are represented by positive operator valued measures (POVMs) and instruments [23, 24, 22]. margin: POVM
@@
“Normality” is a suitable regularity requirement; for a positive map it is equivalent to require the map to be the adjoint of a bounded map on $\mathscr{T}(\mathscr{H})$. Let us recall that an instrument gives the probabilities for the observed values and the conditional state after the measurement. The quantity $\mathcal{I}(\cdot)[\mathds{1}]$ is a POVM, and it gives only the probabilities. Moreover, measurements on a quantum system can be interpreted as involving hybrid systems: a POVM is a channel from a quantum system to a classical one, an instrument is a channel from a quantum system to a hybrid system [43, 2, 7, 8]. When $\mathscr{X}=\mathbb{R}^{s}$ margin: $\mathbb{R}^{s}$ it is usual to take for $\Sigma_{\mathscr{X}}$ the $\sigma$-algebra of the Borel sets, denoted by $\mathcal{B}(\mathbb{R}^{s})$. margin: $\mathcal{B}(\mathbb{R}^{s})$
The classical component shall have values in $\mathbb{R}^{s}$, the classical phase space [7, 13]. The choice of observable and state spaces in the classical case is more involved [7, Sec. 1.1]. To introduce the semigroups associated to the Markov processes on which our classical component will be based, it is convenient to use the space $B_{\rm b}(\mathbb{R}^{s})$ margin: $B_{\rm b}(\mathbb{R}^{s})$ of bounded, Borel measurable complex functions on $\mathbb{R}^{s}$ [38, 37, 39] and its subspace $C_{\rm b}(\mathbb{R}^{s})$ margin: $C_{\rm b}(\mathbb{R}^{s})$ of the continuous bounded functions; both spaces are Banach spaces under the supremum norm [39, p. 6]. We consider complex functions instead of only real ones, because it is more convenient in constructing hybrid systems. Inside a $C^{*}$-algebraic approach, $C_{\rm b}(\mathbb{R}^{s})$ was one of the suggestions for hybrid systems given in [7, p. 7].
@@ 2 Hybrid dynamics and stochastic differential equations
In the quantum theory of measurements in continuous time, many conceptual and technical results were obtained by the quantum trajectory approach, based on stochastic differential equations (SDEs) [22, 31, 21, 23, 32, 24]. By this approach, it is possible to develop a quantum trajectory formulation of a very general class of hybrid dynamics.
Let us start by introducing the Wiener and Poisson processes needed in the construction.
@@
Let $\big{(}\Omega,\mathscr{F},\mathbb{Q}\big{)}$ margin: $\left(\Omega,\mathscr{F},\mathbb{Q}\right)$ be a probability space with a filtration of $\sigma$-algebras $\left\{\mathscr{F}_{t},t\geq 0\right\}$, margin: $\mathscr{F}_{t},\;\mathscr{F}_{t_{-}}$ such that  (a)  $\mathscr{F}=\bigvee_{t\geq 0}\mathscr{F}_{t}$,  (b)  $\mathscr{F}_{t}=\bigcap_{r>t}\mathscr{F}_{r}$,  (c)  $\mathsf{F}\in\mathscr{F},\ \mathbb{Q}(\mathsf{F})=0\ \Rightarrow\ \mathsf{F}\in\mathscr{F}_{0}$; we set also $\mathscr{F}_{t_{-}}=\bigvee_{0\leq r<t}\mathscr{F}_{r}$ [33, pp. 3-4].
Let $W_{k}$, $k=1,\ldots,d$, margin: $W_{k}(t)$ be continuous versions of independent, adapted, standard Wiener processes, with increments independent of the past [34, Sec. I.7].
Let $\mathscr{U}$ be a locally compact Hausdorff space with a topology with a countable basis and $\mathcal{B}(\mathscr{U})$ margin: $(\mathscr{U},\mathcal{B}(\mathscr{U}))$ be the Borel $\sigma$-algebra of $\mathscr{U}$. Let $\nu$ be a Radon measure on $(\mathscr{U},\mathcal{B}(\mathscr{U}))$ with margin: $\nu(\mathrm{d}u)$ margin: $\mathsf{U}_{0}$
@@
Finally, $\Pi(\mathrm{d}u,\mathrm{d}t)$, margin: $\Pi(\mathrm{d}u,\mathrm{d}t)$ is an adapted Poisson point process on $\mathscr{U}\times\mathbb{R}_{+}$ of intensity $\nu(\mathrm{d}u)\mathrm{d}t$; $\Pi$ is independent of the Wiener processes and with increments independent of the past [34, Secs. I.8, I.9].
The simpler case of a finite collection of Poisson processes is presented in Sec. 4.2. Stochastic integrals with respect to the Wiener and Poisson processes are defined, for instance, in [33, 34, 35, 36, 37]; it is also useful to introduce the compensated Poisson measure: margin: $\widetilde{\Pi}(\mathrm{d}u,\mathrm{d}t)$
@@
Moreover, we denote by $\operatorname{\mathbb{E}}_{\mathbb{Q}}[X]$ margin: $\operatorname{\mathbb{E}}_{\mathbb{Q}}$ the mean value of the random variable $X$ with respect to the probability measure $\mathbb{Q}$. In the following we shall make use of Itô’s formula and generalizations; see [34, Sec. II.5], [35, Sec. I.4], [37, Sec. II.7].
@@ 2.1 The classical component
The classical component of the hybrid system is the stochastic process $X(t)$ taking values in the phase space $\mathbb{R}^{s}$ and satisfying the stochastic differential equation (SDE) margin: SDE
$\mathrm{d}X_{i}(t)=c_{i}\big{(}X(t_{-})\big{)}\mathrm{d}t+\sum_{k=1}^{d}a_{i}^{k}\big{(}X(t_{-})\big{)}\mathrm{d}W_{k}(t)\\ {}+\int_{\mathsf{U}_{0}}g_{i}\big{(}X(t_{-}),u\big{)}\widetilde{\Pi}(\mathrm{d}u,\mathrm{d}t)+\int_{\mathscr{U}\setminus\mathsf{U}_{0}}g_{i}\big{(}X(t_{-}),u\big{)}\Pi(\mathrm{d}u,\mathrm{d}t),\qquad i=1,\ldots,s.$ (2)
When considering stochastic processes with diffusion and jumps, it is useful to fix some regularity properties for their trajectories. We always understand to have regular right continuous (RRC)margin: RRC processes, which means processes adapted to the given filtration $\{\mathscr{F}_{t}\}_{t\geq 0}$, with trajectories continuous from the right and with limits from the left [33, Def. 1.5]; indeed, $X(t_{-})$ means the limit from the left.
@@
The $a_{i}^{k}$ and $c_{i}$ are Borel measurable functions from $\mathbb{R}^{s}$ to $\mathbb{R}$; the $g_{i}$ are measurable functions from $\mathbb{R}^{s}\times\mathscr{U}$ into $\mathbb{R}$. We have also, for some $K>0$, and $\forall x,y\in\mathbb{R}^{s}$,
$\left\|c(x)\right\|^{2}+\sum_{k}\left\|a^{k}(x)\right\|^{2}+\int_{\mathsf{U}_{0}}\left\|g(x,u)\right\|^{2}\nu(\mathrm{d}u)\leq K\left(1+\left\|x\right\|^{2}\right),$ (3)
$\left\|c(x)-c(y)\right\|^{2}+\sum_{k}\left\|a^{k}(x)-a^{k}(y)\right\|^{2}+\int_{\mathsf{U}_{0}}\left\|g(x,u)-g(y,u)\right\|^{2}\nu(\mathrm{d}u)\leq K\left\|x-y\right\|^{2}.$ (4)
@@
Under Assumptions 1 and 2, given a random ($\mathscr{F}_{0}$-measurable) initial condition $X(0)$, there is a unique RRC process solving (2); moreover, it turns out to be a time homogeneous Markov process. In particular, $\forall x\in\mathbb{R}^{s}$ there is a unique solution $X^{x}(t)$ margin: $X^{x}(t)$ with deterministic initial condition $X^{x}(0)=x$.
The proof of this Proposition is given in [34, Sec. IV.9]. It is usual to call transition probability (or transition function) [38, 39, 37] the function on $[0,+\infty)\times\mathbb{R}^{s}\times\mathcal{B}(\mathbb{R}^{s})$, defined by
@@
Some properties of the solutions of (2) and of the transition probabilities are collected in Appendix A, where also the associated semigoup is given.
Remark 1.
If for some $i$ we have $c_{i}=0$, $g_{i}=0$, and $a_{i}^{j}=\delta_{ij}$, the corresponding Wiener process $W_{i}$ is among the observed classical processes. If for some $i$ we have $c_{i}=0$, $a_{i}^{\cdot}=0$, and $g_{i}(x,u)=\mathtt{1}_{\mathsf{U}}(u)$, with $\nu(\mathsf{U})<+\infty$, the Poisson process $\Pi(\mathsf{U},(0,t])$ is among the observed classical processes.
The process $X(t)$ is defined under the probability measure $\mathbb{Q}$, as the Wiener processes and the Poisson measure are. However, let us note that $\mathbb{Q}$ will play the role of a reference probability. The physical probability will be constructed in Sec. 2.3.1, in terms of $\mathbb{Q}$ and of the quantum/classical dynamics.

### v2

@@ Abstract
Quantum trajectory techniques have been used in the theory of open systems as a starting point for numerical computations and to describe the monitoring of a quantum system in continuous time. We extend this technique to develop a general approach to the dynamics of quantum/classical hybrid systems. By using two coupled stochastic differential equations, we can describe a classical component and a quantum one which have their own intrinsic dynamics and which interact with each other. A mathematically rigorous construction is given, under the restriction of having a Markovian joint dynamics and of involving only bounded operators on the Hilbert space of the quantum component. An important feature is that, if the interaction allows for a flow of information from the quantum component to the classical one, necessarily the dynamics is dissipative. We show also how this theory is connected to a suitable hybrid dynamical semigroup, which reduces to a quantum dynamical semigroup in the purely quantum case and includes Liouville and Kolmogorov-Fokker-Planck equations in the purely classical case. Moreover, this semigroup allows to compare the proposed stochastic dynamics with various other proposals based on hybrid master equations. Some simple examples are constructed in order to show the variety of physical behaviours which can be described; in particular, a model presenting hidden entanglement is introduced.
Keywords: hybrid dynamics, quantum trajectories, positive operator valued measures, hybrid dynamical semigroup
@@ 1 Introduction
The interest in quantum/classical hybrid systems is old, see for instance [1, 7, 14, 13, 4, 3, 9, 6, 8, 11, 10, 12, 2, 5] and references therein. One of the main motivations of the study of hybrid systems is that the output of a monitored system is classical; then, implicitly or explicitly, the dynamics of quantum/classical systems is involved in the theory of quantum measurements in continuous time and quantum filtering [15, 19, 23, 25, 1, 2, 18, 22, 20, 24, 16, 17, 21, 4, 3, 28, 26, 27]. Moreover, hybrid systems could be used as an approximation to complicate quantum systems, as an effective theory [1, 4, 9, 10, 12, 11, 29]. Another motivation is in the connections between a theory of gravity and quantum theories; perhaps gravity is “classical”, and a dynamical theory of gravity and matter needs a hybrid dynamics [33, 30, 31, 4, 1, 32, 28, 5]. Classical environments could be involved even in revivals of quantum entanglement [34, 35]. A common feature in many of these attempts is that a non-trivial hybrid dynamics is always irreversible, due to the flow of information from the quantum system to the classical one [1, 6, 4, 31, 14, 13, 7, 5, 3].
The main aim of this article is to construct a consistent formulation of the dynamics of quantum/classical hybrid systems, by connecting the quantum trajectory theory, used to describe open and monitored quantum systems [36, 37, 23, 24, 16, 17, 21], to the classical theory of Markov processes and stochastic differential equations [38, 39, 40, 41, 42, 44, 43]. Some proposals of connecting quantum trajectories to hybrid dynamics were given in [5, 28, 3]. The dynamics of the hybrid systems is given in Sec. 2 in terms of coupled stochastic differential equations; only the Markovian case is considered. We also show how this approach is connected to notions appearing in the general formulation of quantum measurements, such as “positive operator valued measures” and “instruments” [24, 22, 23].
In Sec. 3 we show how to associate a suitable semigroup to the constructed dynamics. This hybrid semigroup reduces to a quantum dynamical semigroup [22, 23, 24] in the case of a pure quantum system, while in the pure classical case it includes the Liouville and the Kolmogorov-Fokker-Planck equations [44]. We give also the formal generator of this semigroup, which allows to compare the class of dynamics developed here with the ones obtained by other approaches. This generator also shows analogies and links with the theory of infinitely divisible distributions [44, 45, 46, 13].
@@ 1.1 Some notations
By $\mathscr{S}(\mathscr{H})\subset\mathscr{T}(\mathscr{H})$ margin: $\mathscr{S}(\mathscr{H})$ we denote the set of the statistical operators (the quantum states). Then, general measurements are represented by positive operator valued measures and instruments [23, 24, 22].
@@
“Normality” is a suitable regularity requirement; for a positive map it is equivalent to require the map to be the adjoint of a bounded map on $\mathscr{T}(\mathscr{H})$. Let us recall that an instrument gives the probabilities for the observed values and the conditional state after the measurement. The quantity $\mathcal{I}(\cdot)[\mathds{1}]$ is a positive operator valued measure, and it gives only the probabilities. Moreover, measurements on a quantum system can be interpreted as involving hybrid systems: a positive operator valued measure is a channel from a quantum system to a classical one, an instrument is a channel from a quantum system to a hybrid system [48, 2, 7, 9]. When margin: $\mathbb{R}^{s}$ $\mathscr{X}=\mathbb{R}^{s}$ it is usual to take for $\Sigma_{\mathscr{X}}$ the $\sigma$-algebra of the Borel sets, denoted by $\mathcal{B}(\mathbb{R}^{s})$. margin: $\mathcal{B}(\mathbb{R}^{s})$
We take a classical component with values in $\mathbb{R}^{s}$, the classical phase space [7, 13]. The choice of observable and state spaces in the classical case is more involved [7, Sec. 1.1]. To introduce the semigroups associated to the Markov processes on which our classical component will be based, it is convenient to use the space $B_{\rm b}(\mathbb{R}^{s})$ margin: $B_{\rm b}(\mathbb{R}^{s})$ of bounded, Borel measurable complex functions on $\mathbb{R}^{s}$ [43, 42, 44] and its subspace $C_{\rm b}(\mathbb{R}^{s})$ margin: $C_{\rm b}(\mathbb{R}^{s})$ of the continuous bounded functions; both spaces are Banach spaces under the supremum norm [44, p. 6]. We consider complex functions instead of only real ones, because it is more convenient in constructing hybrid systems. Inside a $C^{*}$-algebraic approach, $C_{\rm b}(\mathbb{R}^{s})$ was one of the suggestions for hybrid systems given in [7, p. 7].
@@ 2 Hybrid dynamics and stochastic differential equations
In the quantum theory of measurements in continuous time, many conceptual and technical results were obtained by the quantum trajectory approach, based on stochastic differential equations (SDEs) [22, 36, 21, 23, 37, 24]. By extending this technique, it is possible to develop a quantum trajectory formulation of a very general class of hybrid dynamics.
The construction is based on two probability measures: a reference probability $\mathbb{Q}$ and a physical probability $\mathbb{P}$. The physical probability will be constructed in Sec. 2.3.1, in terms of $\mathbb{Q}$ and of the quantum/classical dynamics.
In the next Assumption we introduce the reference probability and the Wiener and Poisson processes needed in the construction. They are defined in a filtered probability space satisfying the usual hypothesis, typically assumed in books on stochastic calculus [42, p. 3 and Sec. I.5].
@@
Let $\big{(}\Omega,\mathscr{F},\mathbb{Q}\big{)}$ margin: $\left(\Omega,\mathscr{F},\mathbb{Q}\right)$ be a complete probability space with a filtration of $\sigma$-algebras $\left\{\mathscr{F}_{t},t\geq 0\right\}$, margin: $\mathscr{F}_{t},\;\mathscr{F}_{t_{-}}$ such that  (a)  $\mathscr{F}=\bigvee_{t\geq 0}\mathscr{F}_{t}$,  (b)  $\mathscr{F}_{t}=\bigcap_{r>t}\mathscr{F}_{r}$,  (c)  $\mathsf{F}\in\mathscr{F},\ \mathbb{Q}(\mathsf{F})=0\ \Rightarrow\ \mathsf{F}\in\mathscr{F}_{0}$; we set also $\mathscr{F}_{t_{-}}=\bigvee_{0\leq r<t}\mathscr{F}_{r}$ [38, pp. 3-4], [39, p. 45].
Let $W_{k}$, $k=1,\ldots,d$, margin: $W_{k}(t)$ be continuous versions of independent, adapted, standard Wiener processes, with increments independent of the past [39, Sec. I.7].
Let $\mathscr{U}$ margin: $(\mathscr{U},\mathcal{B}(\mathscr{U}))$ be a locally compact Hausdorff space with a topology with a countable basis and $\mathcal{B}(\mathscr{U})$ be the Borel $\sigma$-algebra of $\mathscr{U}$. Let $\nu$ margin: $\nu(\mathrm{d}u)$, $\mathsf{U}_{0}$ be a Radon measure on $(\mathscr{U},\mathcal{B}(\mathscr{U}))$ with the property:
@@
Finally, $\Pi(\mathrm{d}u,\mathrm{d}t)$ margin: $\Pi(\mathrm{d}u,\mathrm{d}t)$ is an adapted Poisson point process on $\mathscr{U}\times\mathbb{R}_{+}$ of intensity $\nu(\mathrm{d}u)\mathrm{d}t$; $\Pi$ is independent of the Wiener processes and with increments independent of the past [21, Sec. 1], [39, Secs. I.8, I.9].
It is also useful to introduce the margin: $\widetilde{\Pi}(\mathrm{d}u,\mathrm{d}t)$ compensated Poisson measure:
@@
Stochastic integrals with respect to the Wiener and Poisson processes are defined, for instance, in [38, 39, 40, 41, 42]. In particular, the stochastic integrals with respect to $\Pi(\mathrm{d}u,\mathrm{d}t)$ and with respect to $\widetilde{\Pi}(\mathrm{d}u,\mathrm{d}t)$ are defined in [39, Sec. II.3], where also the classes of possible integrands are introduced. In the following we shall make use of Itô’s formula and generalizations; see [39, Sec. II.5], [40, Sec. I.4], [42, Sec. II.7]. The simpler case of a finite collection of Poisson processes is presented in Sec. 4.2.
We shall denote by $\operatorname{\mathbb{E}}_{\mathbb{Q}}[X]$ margin: $\operatorname{\mathbb{E}}_{\mathbb{Q}}$ the mean value of the random variable $X$ with respect to the probability measure $\mathbb{Q}$. As a real random variable is a measurable function from $\Omega$ to $\mathbb{R}$ we can write $X:\omega\in\Omega\to X(\omega)\in\mathbb{R}$; then, the mean value can be written as $\operatorname{\mathbb{E}}_{\mathbb{Q}}[X]=\int_{\Omega}X(\omega)\mathbb{Q}(\mathrm{d}\omega)$.
@@ 2.1 The classical component
The classical component of the hybrid system is the stochastic process $X(t)$ taking values in the phase space $\mathbb{R}^{s}$ and satisfying the stochastic differential equation
$\mathrm{d}X_{i}(t)=c_{i}\big{(}X(t_{-})\big{)}\mathrm{d}t+\sum_{k=1}^{d}b_{i}^{k}\big{(}X(t_{-})\big{)}\mathrm{d}W_{k}(t)+\int_{\mathsf{U}_{0}}g_{i}\big{(}X(t_{-}),u\big{)}\widetilde{\Pi}(\mathrm{d}u,\mathrm{d}t)\\ {}+\int_{\mathscr{U}\setminus\mathsf{U}_{0}}g_{i}\big{(}X(t_{-}),u\big{)}\Pi(\mathrm{d}u,\mathrm{d}t),\qquad i=1,\ldots,s.$ (2)
When considering stochastic processes with diffusion and jumps, it is useful to fix some regularity properties for their trajectories. We always understand to have regular right continuous processes, which means processes adapted to the given filtration $\{\mathscr{F}_{t}\}_{t\geq 0}$, with trajectories continuous from the right and with limits from the left [38, Def. 1.5]; indeed, $X(t_{-})$ means the limit from the left.
@@
The $b^{k}$ and $c$ are Borel measurable functions from $\mathbb{R}^{s}$ to $\mathbb{R}^{s}$; the $g$ are measurable functions from $\mathbb{R}^{s}\times\mathscr{U}$ into $\mathbb{R}^{s}$. We have also, for some $K>0$, and $\forall x,y\in\mathbb{R}^{s}$,
$\left\|c(x)\right\|^{2}+\sum_{k}\left\|b^{k}(x)\right\|^{2}+\int_{\mathsf{U}_{0}}\left\|g(x,u)\right\|^{2}\nu(\mathrm{d}u)\leq K\left(1+\left\|x\right\|^{2}\right),$
$\left\|c(x)-c(y)\right\|^{2}+\sum_{k}\left\|b^{k}(x)-b^{k}(y)\right\|^{2}+\int_{\mathsf{U}_{0}}\left\|g(x,u)-g(y,u)\right\|^{2}\nu(\mathrm{d}u)\leq K\left\|x-y\right\|^{2}.$
@@
Under Assumptions 1 and 2, given a random ($\mathscr{F}_{0}$-measurable) initial condition $X(0)$, there is a unique regular right continuous process solving (2); moreover, it turns out to be a time-homogeneous Markov process. In particular, $\forall x\in\mathbb{R}^{s}$ there is a unique solution $X^{x}(t)$ margin: $X^{x}(t)$ with deterministic initial condition $X^{x}(0)=x$.
The proof of this Proposition is given in [39, Sec. IV.9]; see, in particular, Theorem 9.1 of [39]. Note that no condition is given on $g(x,u)$ for $u\in\mathscr{U}\setminus\mathsf{U}_{0}$, apart from being jointly measurable in $(x,u)$. This is due to the fact that both $\Pi$ and $\nu$ are finite measures on $\mathscr{U}\setminus\mathsf{U}_{0}$.
It is usual to call transition probability (or transition function) [43, 44, 42] the function on $[0,+\infty)\times\mathbb{R}^{s}\times\mathcal{B}(\mathbb{R}^{s})$, defined by
@@
Some properties of the solutions of (2) and of the transition probabilities are collected in A, where also the associated semigoup is given.
If for some $i$ we have $c_{i}=0$, $g_{i}=0$, and $b_{i}^{j}=\delta_{ij}$, the corresponding Wiener process $W_{i}$ is among the observed classical processes. If for some $i$ we have $c_{i}=0$, $b_{i}^{\cdot}=0$, and $g_{i}(x,u)=\mathtt{1}_{\mathsf{U}}(u)$, with $\nu(\mathsf{U})<+\infty$, the Poisson process $\Pi(\mathsf{U},(0,t])$ is among the observed classical processes.